Determine complex number(s) z such that z- i , iz- i and z-iz have the same modulus


1
Expert's answer
2021-11-10T10:13:03-0500

"|z-i|=|iz-i|=|z-iz|"

"z=x+iy"

"z-i=x+i(y-1)"

"iz-i=-y+i(x-1)"

"z-iz=(x+y)+i(y-x)"


"|z-i|=\sqrt{x^2+(y-1)^2}"

"|iz-i|=\sqrt{y^2+(x-1)^2}"

"|z-iz|=\sqrt{(x+y)^2+(y-x)^2}"


"x^2+(y-1)^2=y^2+(x-1)^2"

"-2y=-2x"

"x=y"


"y^2+(x-1)^2=(x+y)^2+(y-x)^2"

"2x^2-2x+1=2x^2"

"x=y=1\/2"


"z=0.5+0.5i"